Skip to Main Content
PTAGIS logo link back to home page

Interrogation Site Metadata

Search by typing either a site name, site code, organization, or site type.

ICH - Ice Harbor Dam (Combined)

Status:   Active - Non-Operational
Site Type: Combined Dam Location  |  Organization: Pacific States Marine Fisheries Commission
General Operational Period: Year Round
Start Date: 4/8/2005 9:00 AM

Ice Harbor Dam Adult Fishways (both) and Full Flow Bypass

Current Operations

Last File Loaded: ICH-2021-297-P-008.xml
Date Last File Loaded: 10/24/2021 10:13 AM
Date Last File Opened: 10/24/2021 7:08 AM
Date Last File Closed: 10/24/2021 8:08 AM
Last Observation Tag: 3DD.003D829819
Last Observation Date: 10/23/2021 4:52 PM
Detection Summary: Last Day's Observations

Contacts

Primary Contact:

Don Warf
dlwarf@psmfc.org
509-735-2773

Data Steward:

Nicole Tancreto
ntancreto@psmfc.org
503-595-3100

Location

Coordinates: 46.249527, -118.879964  |  River Km: 522.016

Operational History

Operational history describes the period the site was interrogating for PIT tags. The comments provide anecdotal information about the operation of the site during this period. Site operational periods are updated typically at the end of the calendar year.

Year Start Date End Date Comments
2021
2020 1/1/2020 12:00 AM 11/1/2021 12:00 AM Full flow operated March 23 - November 23. Left ladder was dewatered January 7-31. Right ladder was dewatered February 4 - March 2.
2019 1/1/2019 12:00 AM 1/1/2020 12:00 AM Left ladder dewatered January 2 - February 28. Full flow watered up March 21 - December 19.
2018 1/1/2018 12:00 AM 1/1/2019 12:00 AM Right ladder was dewatered from Jan 5 - Feb 1. Left ladder was dewatered from Feb 1 - 28. Full flow watered up March 22 - December 21.
2017 1/1/2017 12:00 AM 1/1/2018 12:00 AM Right ladder dewatered Jan 3-27, left ladder dewatered Jan 30-Feb 28. Full flow watered up March 22-Dec 20.
2016 1/1/2016 12:00 AM 1/1/2017 12:00 AM Left ladder dewatered January 4-February 4. Right ladder dewatered February 8-29. Full flow operational from March 24-December 21.
2015 1/1/2015 12:00 AM 1/1/2016 12:00 AM Left ladder was dewatered January 5-26. Right ladder was dewatered January 26-Febarury 24. Juvenile bypass operational March 18. Juvenile bypass operated March 18 - December 18.
2014 1/1/2014 12:00 AM 1/1/2015 12:00 AM Left ladder was dewatered January 6-February 3. Right ladder was dewatered February 10-24. Juvenile bypass operationl March 17-December 12.
2013 1/1/2013 12:00 AM 1/1/2014 12:00 AM Right ladder was dewatered from January 2-8. Left ladder was dewatered from January 14-March 5. The full flow juvenile bypass watered up on March 12 and was dewatered for the season on December 23.
2012 1/1/2012 12:00 AM 1/1/2013 12:00 AM The Left (South) ladder was dewatered from January 1-26. The right (north) ladder was dewatered from January 28 - February 21. The full flow watered up on March 19 and was dewatered on December 19.
2011 1/1/2011 12:00 AM 1/1/2012 12:00 AM The IHR North adult fishway was dewatered between Jan. 3-14 for annual maintenance. The South ladder was dewatered from Jan. 18 until Mar. 2 for annual maintenance. Full flow watered up on March 23 and was dewatered for the season on December 21.
2010 1/1/2010 12:00 AM 1/1/2011 12:00 AM The right (North) ladder was dewatered between January 4-13, 2010. The left (South) ladder was dewatered between Jan. 19 and Feb. 23. The juvenile bypass watered up on March 8, and was dewatered on Dec. 20.
2009 1/1/2009 12:00 AM 1/1/2010 12:00 AM The right (North) ladder was dewatered between January 6-21, 2009. The juvenile fish bypass was dewatered between January 12 and March 15. The left (South) ladder was dewatered between January 26 - February 25. The juvenile fish bypass (full-flow) pipe was dewatered on December 17.
2008 1/1/2008 12:00 AM 1/1/2009 12:00 AM The left (South) ladder was dewatered from approximately 10AM on Jan. 9 until 10AM on Jan. 19. The right (North) ladder was dewatered for annual maintenance from 9AM on Feb. 4 until 2PM on Feb. 13. The full-flow juvenile fish bypass watered-up for the season on March 17.
2007 1/1/2007 12:00 AM 1/1/2008 12:00 AM The North ladder was out of service from Jan. 3 through Jan. 9 for its scheduled annual inspection and maintenance. The South Ladder was similarly dewatered between Jan. 17 and Feb. 14. The fish bypass watered-up on March 5 and remained in service until December 18.
2006 1/1/2006 12:00 AM 1/1/2007 12:00 AM The South ladder was dewatered for annual maintenance on January 3, 2006; it was returned to service by Jan. 18. The full flow pipe for juvenile fish bypass watered up on March 2, and dewatered on December 20.
2005 4/8/2005 9:00 AM 1/1/2006 12:00 AM This site definition is new in 2005. It replaces the previous IHA definition with the addition of detectors in the full flow section of the juvenile bypass system, and incorporates the existing detectors in the North and South shore adult fish ladders. Per normal operating procedures, the Full Flow Bypass pipe was dewatered at ~9:00 AM on December 12, 2005, and will remain OOS through the winter.

Configuration

Configuration metadata includes details about the equipment used at the site to detect and record PIT tags. A new configuration sequence record is created when equipment is added or removed from the site and the period of time it was active. A configuration sequence without an End Date is currently active.

Select Configuration Number:
Start Date:  12/9/2015
End Date:  ----

Site Configuration Diagrams:

Equipment History

Select a type to view the history of equipment installed at this site. A missing End Date implies the information about the equipment is current.

Select Equipment Type:
Equipment Code Description Start Date End Date

Photos

Photos are contributed from stewards of the site.

Event Log Summary

This displays the last 25 events, to see the full history, log in to the Dashboard and go to Interrogation Site > Event Logs.

Event Date Event Type Comments
10/23/2021 8:56:51 AM ErrorsCleared EVents are likely due to noise events 
10/22/2021 7:32:12 AM ErrorsCleared The noise is being investigated by COE personnel.
10/21/2021 7:38:19 AM ErrorsCleared Noise on these systems has been identified and is being investigated by COE personnel.
10/20/2021 7:43:38 AM ErrorsCleared
10/17/2021 11:02:39 AM ErrorsCleared The noise has been linked to new VFDs installed to run the lock gates. Will work with the COE to try to resolve it. 
10/16/2021 10:48:43 AM ErrorsCleared The noise has been linked to new VFDs installed to run the lock gates. Will work with the COE to try to resolve it. 
10/16/2021 10:45:12 AM ErrorsCleared Re-tune transceiver on next visit.
10/14/2021 7:31:25 AM ErrorsCleared 0A will need to manually tuned during the next site visit. The noise event is well documented.
10/13/2021 8:45:21 AM ErrorsCleared The noise has been linked to new VFDs installed to run the lock gates. Will work with the COE to try to resolve it. 
10/13/2021 8:44:51 AM ErrorsCleared Re-tune transceiver on next visit.
10/12/2021 9:16:45 AM ErrorsCleared The noise has been linked to new VFDs installed to run the lock gates. Will work with the COE to try to resolve it. 
10/12/2021 9:15:15 AM ErrorsCleared Re-tune transceiver on next visit.
10/12/2021 8:56:28 AM ErrorsCleared Re-tune transceiver on next visit.
10/11/2021 9:35:22 AM ErrorsCleared 0A will need to manually tuned during the next site visit. The noise event is well documented.
10/10/2021 8:00:50 AM ErrorsCleared 0A will need to manually tuned during the next site visit. The noise event is well documented.
10/8/2021 7:11:30 AM ErrorsCleared 0A will need to manually tuned during the next site visit. The noise event is well documented.
10/7/2021 7:10:50 AM ErrorsCleared  Noise alarms generated by Lock gate operations. The tuning on 0A will be done during the next site visit.
10/6/2021 7:28:14 AM ErrorsCleared Noise alarms generated by Lock gate operations. The tuning on 0A will be done during the next site visit.
10/5/2021 7:21:05 AM ErrorsCleared To be investigated.
10/3/2021 8:42:47 AM ErrorsCleared The noise has been linked to new VFDs installed to run the lock gates. Will work with the COE to try to resolve it. 
10/3/2021 8:42:34 AM ErrorsCleared Re-tune transceiver on next visit.
10/2/2021 8:12:29 AM ErrorsCleared The noise has been linked to new VFDs installed to run the lock gates. Will work with the COE to try to resolve it. 
9/30/2021 9:14:05 AM ErrorsCleared The noise has been linked to new VFDs installed to run the lock gates. Will work with the COE to try to resolve it. 
9/30/2021 9:13:36 AM ErrorsCleared The system experienced a single noise spike for a short duration, and is otherwise quiet. The readers ability to detect PIT tags and timer tags were not impacted by the noise.
9/28/2021 10:16:30 AM ErrorsCleared

Archived Event Logs